Reports 1-1 of 1 Clear search Modify search
shinji.miyoki - 17:09 Wednesday 21 April 2021 (16532) Print this report
Comment to Setting the mirror to nominal height (16521)

I found the temp increase from 20:00 on 20th around 16:00 on 21st after Fabian's notice. I suspected the stop of SR-C cooler. Yoshimura-san found it actually stopped, then he restarted it. The stop reason was same as previous trouble.

Search Help